OpenStudy (anonymous):

sin2xcosx-cos2xsinx=-1/2 I get its sin(a-b) I just don't know how to apply it

OpenStudy (psymon):

Well, it looks like that a would be 2x and b would be x in this case. So this is just a difference of sines formula originating from just sin(x). So looks like all you have to do is solve for sin(x) = -1/2

OpenStudy (anonymous):

i need to answer for A= and b = though ... i feel like im just overthinking the problem

OpenStudy (psymon):

Well a is just 2x and b is just x. Unless you mean that we need to find sin2x= -1/2 and sinx = -1/2. Otherwise I think you're justoverthinking it.

OpenStudy (anonymous):

so A would just be 2B essentially ?

OpenStudy (psymon):

Right. Because the difference of sines is sinAcosB - cosAsinB. So if all you need are A and B, then you just pluck them right from the equation, making A = 2x and B = x. Now if you had to actually solve for x, then you have (sin2x - x) = -1/2: sinx = -1/2

OpenStudy (anonymous):

ok thank you very much

OpenStudy (psymon):

Mhm :3